OpenStudy (anonymous):

could someone please help explain how find the exact values would appreciate it. Ex: cos 135 degrees

OpenStudy (anonymous):



OpenStudy (anonymous):

Hello, are you familiar with the unit circle?

OpenStudy (anonymous):

For example, \[\cos (30)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

hey and uhm no not that much but I know what it is. I know the values and etc like you labeled cos 30

OpenStudy (anonymous):

Okay, what you really want to do is checkout the unit circle, at each major point on the circle is an x,y coordinate

OpenStudy (anonymous):

sin(x) is equivalent to the y value and cos(x) is equivalent to the x value

OpenStudy (anonymous):

okay so suppose the angle here I labeled as an example is 135 but what if its like above 360 for example 460 degrees?

OpenStudy (anonymous):

the easiest way is to convert to radians and subtract 2 pi

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

Would mean alot if you could help just began this and already confused and got a test tomorrow.

OpenStudy (anonymous):

I know how to convert to radians not the rest.

OpenStudy (anonymous):

on the unit circle, take a look at 135 deg

OpenStudy (anonymous):

cos(x) is equivalent of the x value, you see \[-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

that is the exact value

OpenStudy (anonymous):

yes x is is that in the pair of coordinates

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

could you please do me a favor

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I did this on my own my study guide please verify or criticize on my wrong doings, I was to once again find the exact value for an angle of 125 degrees so I made a graph and did 180-125 and got 55 and since tan 1255 is in the second quadrant and negative i got the final answer of -tan 55 degrees. would that be correct?

OpenStudy (anonymous):

tan 125* not 1255 sorry

OpenStudy (anonymous):

Yes, that is

OpenStudy (anonymous):

phew! thanks. but one more question could you please help me with an example if the angle was greater than 360 degrees?

OpenStudy (anonymous):

Whats your question?

OpenStudy (anonymous):

cos 850 degrees how would I solve that for the exact value?

OpenStudy (anonymous):

main point or question being what to do if the value is greater than 360

OpenStudy (anonymous):

Round to the nearest tenth. Well, that is if your dong trigonometry. Are you dong Trigonometry?

OpenStudy (anonymous):

yes I am.

OpenStudy (anonymous):

Alright, well anything that is equal to 360 degrees is a circle. But, cos85 = -36.8292. What you do next is that you round -36.8292 to the nearest 10th.

OpenStudy (anonymous):

uhm I wrote 850 but we still round?

OpenStudy (anonymous):

You have to solve cos 850 to get your answer..

OpenStudy (anonymous):

alright I think I got the idea thanks :)

OpenStudy (anonymous):

No problem.